# Spectral Bottleneck in Deep Neural Networks: Noise is All You Need

### v1

@@ 5.3 Image denoising
The robustness of different INR architectures is assessed for the canonical image denoising. Firstly, a clean signal $f(\mathbf{x})$ is corrupted by additive white Gaussian noise $\eta(\mathbf{x})\sim\mathcal{N}(0,\sigma^{2})$ such that $\tilde{f}(\mathbf{x})=f(\mathbf{x})+\eta(\mathbf{x})$ achieves a signal-to-noise ratio of $\mathrm{SNR}=5\ \mathrm{dB}$. The task is to reconstruct $f$ from $\tilde{f}$. We adopt an unsupervised denoising strategy similar to Noise2Self batson2019noise2self , training the INR directly on the noisy input while reserving a small subset of pixels for ‘J-invariant’ validation. A predictor is J-invariant if, for any pixel $i$, the prediction at $i$ does not depend on the noisy value at $i$ itself. This restriction forces the model to reconstruct structure from spatial context rather than memorize pixelwise noise. The held-out set then provides an unsupervised early-stopping criterion that prevents overfitting to noise while retaining underlying structure. Ground-truth images are used strictly for evaluation (e.g., PSNR, SSIM) and never for model selection. For the image quality evaluation, along with PSNR we report structural similarity wang2004image , mean absolute error (MAE), LPIPS zhang2018unreasonable , and DIST ding2020image . For clarity, arrows ($\uparrow$ / $\downarrow$) are used in tables and figures to indicate whether higher or lower values correspond to better performance.
@@
As shown in Fig. 11, SIREN and FINER oversmooth fine-scale details, while WIRE preserves global appearance but leaves residual noise. SIREN2 achieves the sharpest and most faithful reconstructions (with the best SSIM and LPIPS), suppressing noise while retaining textures and edges.
@@
5.4 Audio denoising
Table 3: Audio denoising. Best PSNR ($\uparrow$) in dB for different audio clips using FINER++ and SIREN2.
@@ 5.4 Audio denoising
We adopt the same Noise2Self-inspired DIP training procedure from Sec. 5.3 for the present audio denoising experiments. Ground-truth audio signal is used only for evaluation. While the general denoising framework is identical, audio signals present distinct challenges. Unlike natural images, which concentrate most energy in low frequencies, audio signals often exhibit a relatively broadband structure (e.g., higher harmonics in music, ambient noise), causing stronger overlap between the underlying signal and Gaussian noise. This overlap makes simple frequency-selective filtering less effective. In an a-priori setting, where no ground-truth information is available, INR-based denoising can provide a useful alternative.
in Sec. 5.1 on supervised audio fitting task, $\omega_{0}$ was scaled by a factor of 100 consistently for all network architectures. However, for the present audio-denoising task, a large first-layer periodicity consistently degraded denoising for SIREN, Gauss, WIRE, and FINER, with best denoising $\mathrm{PSNR}<25\,\mathrm{dB}$ for different signals; due to the presence of broadband noise in the noisy signal. We therefore avoid first-layer $\omega_{0}$ scaling for denoising and report results for FINER++ and SIREN2 architectures only which does not use any first layer $\omega_{0}$ scaling. Table 3 and Fig. 12 report the accuracy of denoised reconstructions for bach.wav audio clip after 20,000 epochs for FINER++ and SIREN2. The noise scales $s_{0}$ and $s_{1}$ in SIREN2 provide controllable filtering, allowing a principled adjustment of frequency support to match the noise characteristics. For FINER++, varying the bias ranges was evaluated. The results in Table 3 show that both networks achieve competitive accuracy in the recovery of the underlying signal.
@@
Figure 12: Audio denoising. Log-spectrogram comparison for the Bach audio denoising task. The ground-truth signal (top left) is corrupted with additive Gaussian noise at $5\,\text{dB}$ SNR. Reconstructions using SIREN2 recover the smooth structures of the signal effectively.
6 Conclusions
This work identifies and addresses a fundamental limitation of implicit neural representation networks. We show that sinusoidal representation networks (SIRENs) sitzmann2020implicit when not initialized appropriately are prone to fail at fitting high-frequency dominant signals. In extreme cases when the frequency support of the network does not align the frequency spectrum of the target, we identify a spectral bottleneck phenomenon, where the network yields a zero-valued output, failing to capture even those components of the target signal that are within its representational capacity. To address this issue, we propose a target-aware Gaussian weight perturbation scheme WINNER. The proposed scheme introduces Gaussian perturbations to the uniformly distributed weights of a base SIREN at initialization. These perturbations in turn affect the pre-activation distributions of each network layer and their spectra. Similar to random Fourier embeddings tancik2020fourier , this weight perturbation scheme can be used to control the empirical NTK and its eigenbasis, with the benefit of not introducing additional trainable parameters. We achieve state-of-the-art results in audio fitting and demonstrate notable improvements over existing methods in image fitting and denoising tasks. Beyond signal reconstruction, our approach suggests new avenues for adaptive initialization strategies across computer vision tasks that require fine-scale detail.
@@ 6 Conclusions
Limitations. (a) The proposed initialization scheme depends on the prior knowledge of the target to compute the perturbation scales $s_{0}$ and $s_{1}$, making it not applicable directly for cases where the target is unknown a priori (such as denoising or solving a initial/boundary value problem governed by a PDE). This could be addressed by estimating the scales from representative samples or adapting them during early training. The same limitation also applies to Random Fourier embeddings. (b) The proposed SIREN2, along with the other architectures examined in this study, are highly sensitive to the learning rate and it’s decay schedule. Without a scheduler, training often results in a strongly oscillatory PSNR evolution resulting in poor convergence. We recommend a decay rate of 1-2% every 20 epochs with an initial learning rate of $10^{-4}$ for both audio and image fitting to achieve stable, non-oscillatory PSNR evolution.
Broader impact. Improved signal representation through INRs can benefit diverse AI applications such as super-resolution kim2022learning , inpainting kim2022learning , denoising lanzendorfer2023siamese ; fan2019brief , speech synthesis, audio event detection, compression, and captioning. Beyond audio, they are well suited for modeling other 1D waveforms including sensor time-series (e.g., accelerometer, temperature, ECG, EEG), financial time-series, seismic signals, and electronic measurements. Moreover, since INRs provide direct access to spatio-temporal derivatives via automatic differentiation, they can also be applied to solving differential equation-based forward and inverse problems in a mesh-free setting chen2023implicit ; akhare2025implicit ; song2022versatile . The proposed SIREN2 further extends to computer vision tasks such as audio-video fitting, 3D shape representation, and NeRF-style scene reconstruction mildenhall2021nerf .
@@ Acknowledgments
The authors HC, SP, DVS, and IZ gratefully acknowledge the financial assistance provided by Technion - Israel Institute of Technology during the course of present work. HC thanks Sanketh Vedula (Princeton University) for directing him to the work of Sitzmann et al. sitzmann2020implicit ; experimenting with it helped the development of noise scheme presented in this work.
Data and code availability
The data and Python implementation of the experiments presented in this work are made publicly available at https://github.com/hemanthgrylls/SIREN-square.git
Supplementary Information
Appendix A Activation Distribution and Spectra of SIREN and SIREN2
In Section 4, we showed that adding noise to uniformly distributed weights can alter the spectral profile and NTK characteristics of a SIREN network. This enables the design of target-aware initializations for improved learning efficiency. Here we empirically show how these perturbations affect both the distributional properties of activations and their spectral content across layers, providing evidence that SIREN2 modifies the flow of frequency information through the network while largely preserving the favorable statistical structure of the baseline SIREN.
Distributions and Spectra at initialization (Fig. 13). To assess the statistical properties of the proposed weight perturbation scheme, we empirically evaluate a four-layer SIREN (baseline) and SIREN2 networks, each with 2048 hidden features and 16,384 inputs uniformly sampled in [1,1][-1,1]. The distributions and power spectral densities summarized in Fig. 13 show that SIREN2 largely preserves the distributional shape of SIREN, except for the layer-1 pre-activations, which appears to follow a Laplace distribution. Noise addition with WINNER (SIREN2) is observed to alter the spectral profile of the distributions, with a consistent increase in high-frequency power across all layers and a notable reduction in low-frequency power for the layer-3 and layer-4 distributions. Such an change of spectra across the network layers helps mitigate the vanishing gradients problem noted in Fig. 4 when fitting signals dominated by very high frequencies.
Distributions and Spectra at the start and end of training (Fig. 14). We further analyze SIREN and SIREN2 at initialization and after $10^{4}$ iterations when fitting the high-frequency dominant target tetris.wav (150,000 samples) using a four-layer, 256-hidden-unit architecture. tetris.wav was chosen for it’s high-frequency nature. For SIREN, the spectra of distributions remain unaltered throughout the training. In contrast, the spectra of SIREN2 downstream of layer-3 evolve during the training, with high-frequency energy persisting in deeper layers and aligning more closely with the target spectrum. The increase in the standard deviation of distribution downstream of layer-3 can also be acknowledged.
Figure 13: WINNER initialization (used in SIREN2) increases spectral energy in the high-frequency range. Pre- and post-activation distributions and their power spectral densities at initialization for a four layer SIREN (a) and SIREN2 (b) networks. The thick black dashed lines show analytical estimates from the present Theorem 3.1 and Sitzmann et al. [1], closely matching empirical data. In SIREN2, the influence of Gaussian noise (of scales $s_{0}$ and $s_{1}$) can be seen on the spectra pre- and post-activation in the rightmost column. The circled regions with arrows highlight the reduction of spectral energy at low frequencies and the corresponding increase at high frequencies.
Figure 14: Pre-activation statistics of SIREN and $\mathbf{SIREN^{2}}$ at the start and end of training. Distributions and cumulative power spectra of hidden-layer pre-activations and outputs at epoch 0 and $10^{4}$ when fitting tetris.wav with a four-layer, 256-width network. $\mathrm{SIREN}^{2}$ maintains broader high-frequency support during training, enabling recovery of the high-frequency target, whereas SIREN exhibits a rapid loss of high-frequency components across layers.
Appendix B Additional Experiments
B.1 Reducing Parameter Count using Tensor Train networks
We investigate Tensor Train (TT) factorization to reduce the parameter count without degrading reconstruction fidelity. In the four-layer SIREN2 configuration shown in Fig. 15, we replace the fourth dense hidden layer with a low-rank TT linear layer implemented using torchtt [57]. The TT parameterization constrains the weight tensor by prescribed ranks, which lowers storage and compute relative to a dense layer while retaining expressive capacity when ranks are chosen appropriately. Trained for 10,00010{,}000 epochs under the same settings as the dense baseline, the TT-based model attains higher PSNR with fewer parameters. Because TT factorization is complementary to post-training compression, we expect additional savings when combined with quantization techniques [32]. A comprehensive study of rank selection, the accuracy versus compression trade-off, and the interaction with quantization is left to future work.
Figure 15: Four-layer SIREN2 architecture with the fourth dense hidden layer replaced by a tensor train ($T$) linear map that factorizes the $300\times 300$ weight into cores $G_{1}\!\in\!\mathbb{R}^{1\times 10\times 8}$, $G_{2}\!\in\!\mathbb{R}^{8\times 30\times 8}$, and $G_{3}\!\in\!\mathbb{R}^{8\times 30\times 1}$ with $T$ rank $r_{k}=8$. Weights follow the labeled SIREN2 initialization; the TT parameterization reduces parameters and compute while preserving reconstruction fidelity.
Table 4: Performance of tensor train networks for audio reconstruction.
SIREN2
Dense $T$
[…]
Total parameters 198145 182426 (8% $\downarrow$)
Peak PSNR (dB) ($\uparrow$):
relay.wav 71.73 74.35
bach.wav 58.74 61.4
whoosh.wav 44.84 51.95
B.2 Image Fitting on Kodak and DTD datasets
Table 5 reports the average PSNR values for images from the Kodak [45] dataset and for two texture classes from the DTD [46] dataset, namely braided and woven, in RGB mode. The results show that SIREN2 consistently outperforms the baseline SIREN, similar to the trend observed in the grayscale results presented in Table 2. All experiments were conducted using a four-layer network with 256 features.
Table 5: Average reconstruction PSNR ($\uparrow$) of various networks in fitting DTD [46] and Kodak [45] datasets in RGB mode. WIRE [40], ReLU-PE, and Gauss [42] are not presented as they perform relatively poorly compared to SIREN2 and FINER.
# images SIREN SIREN2 (present) FINER
DTD braided dataset 120 45.75 47.15 47.93
DTD woven dataset 120 39.08 40.95 41.64
Kodak dataset 24 35.89 37.48 37.50
B.3 Image Denoising
Following the setup detailed in Sec.5.3, we further evaluate the performance of SIREN2 on image denoising tasks. A direct comparison is made against the original SIREN, and the latest architectures FINER [19], and WIRE [40]. As shown in Figure 16, our model consistently demonstrates superior performance for both the images tested.
Refer to caption(a)
Refer to caption(b)
Figure 16: Denoising results for two images using different network architectures. Top: Sparrow image. Bottom: Galaxy image. The insets highlight regions of interest, illustrating the improved preservation of fine details and overall denoising performance of SIREN2.
B.4 Fitting 3D shapes
We train implicit neural network to fit signed distance functions (SDFs) to oriented pointclounds, following the general implicit representation framework of Sitzmann et al.[1], but with a composite loss formulation closer to recent SDF reconstruction methods [58, 59]. For each batch, 3D points $(\mathbf{x})$ are sampled both near the surface and in the surrounding volume. The network outputs signed distances and surface normals, optimized with following losses
  • SDF loss for consitency
  • Eikonal loss to enforce unit gradient norm
  • Normal loss to align predicted and reference normals
  • Far/outside losses to discourage suprious zero-crossings away from the surfaces
Both FINER++ and SIREN2 were trained under identical conditions for 10,000 epochs. Reconstructions reached peak PSNR values of $51$ dB for FINER++ and $55$ dB for SIREN2. To assess geometric fidelity, we visualize the distance between reconstructed surfaces and the ground-truth mesh. In Fig. 17, vertex-wise errors are color-coded (blue: low error, red: high error). Compared to FINER++, SIREN2 yields visibly lower reconstruction error, especially in regions with fine-scale curvature and high-frequency detail. While FINER++ oversmooths sharp features and introduces local distortions, SIREN2 better preserves geometric structure, leading to tighter alignment with the ground truth.
Figure 17: Comparison of FINER++ and SIREN2 for fitting a signed distance function (SDF) from an oriented point cloud. Both methods produce high-fidelity reconstructions, with SIREN2 performing marginally better. Other methods namely baseline SIREN, WIRE, and FINER performed even poor in comparison; as a results they are not presented.
Appendix C Summary of Computed Spectral Centroids and Noise Scales
Table 6 reports the spectral centroid values $\psi$ together with the corresponding noise scales $s_{0}$ and $s_{1}$, computed using Eqn. 11, for both audio and image targets considered in Sec. 5.1 and Sec. 10. For audio signals, we consider only the first 150,000150{,}000 samples and normalize amplitudes to the range [1,1][-1,1] to ensure consistency across files.
Rationale behind using different noise scales for different targets.
A single choice of noise scales is inadequate for all signals, since the frequency content of the target fundamentally determines the effective frequency support required of the network. Targets with dominant high-frequency components demand broader frequency support, whereas smoother signals require comparatively narrower support (meaning smaller noise scales). To account for this, the proposed scheme computes $s_{0}$ and $s_{1}$ adaptively from the spectral centroid $\psi$ and the channel dimension of the input. While the present formulation is deliberately simple and does not yet yield optimal reconstruction accuracy, it demonstrates the need for a principled adaptation mechanism. A deeper analysis is required to rigorously establish the relationship between spectral descriptors of the target and the optimal choice of noise scales. In particular, it is unlikely that the spectral centroid $\psi$ alone is sufficient; additional statistics of the spectrum are expected to play a critical role in determining the optimal noise scales.
Table 6: Spectral Centroid ($\psi$) and the respective noise scales $s_{0}$, $s_{1}$] computed using Eqn. 11 for audio and image fitting tasks in Sec. 5.1 and 10.
File name Task $\psi$ $s_{0}$ $s_{1}$
tetris.wav Audio fitting 0.5732 3436 1.72
tap.wav Audio fitting 0.7264 3478 2.18
whoosh.wav Audio fitting 0.5266 3412 1.58
radiation.wav Audio fitting 0.8368 3489 2.51
arch.wav Audio fitting 0.4996 3394 1.5
relay.wav Audio fitting 0.6310 3457 1.89
voltage.wav Audio fitting 0.4540 3354 1.36
foley.wav Audio fitting 0.1772 2487 0.53
shattered.wav Audio fitting 0.3942 3278 1.18
bach.wav Audio fitting 0.0737 1410 0.22
birds.wav Audio fitting 0.1789 2499 0.54
noise.png Image fitting 0.5934 47 0.24
camera.png Image fitting 0.3121 39 0.12
castle.jpg Image fitting 0.1097 21 0.04
rock.jpg Image fitting 0.4055 43 0.16
Dragon 3D shape fitting 0.0642 25 1e-3

### v2

@@ VI-C Image Denoising
The robustness of different INR architectures is assessed for the canonical image denoising. Firstly, a clean signal $f(\mathbf{x})$ is corrupted by additive white Gaussian noise $\eta(\mathbf{x})\sim\mathcal{N}(0,\sigma^{2})$ such that $\tilde{f}(\mathbf{x})=f(\mathbf{x})+\eta(\mathbf{x})$ achieves a signal-to-noise ratio of $\mathrm{SNR}=5\ \mathrm{dB}$. The task is to reconstruct $f$ from $\tilde{f}$. We adopt an unsupervised denoising strategy similar to Noise2Self batson2019noise2self, training the INR directly on the noisy input while reserving a small subset of pixels for ‘J-invariant’ validation. A predictor is J-invariant if, for any pixel $i$, the prediction at $i$ does not depend on the noisy value at $i$ itself. This restriction forces the model to reconstruct structure from spatial context rather than memorize pixelwise noise. The held-out set then provides an unsupervised early-stopping criterion that prevents overfitting to noise while retaining underlying structure. Ground-truth images are used strictly for evaluation (e.g., PSNR, SSIM) and never for model selection. For the image quality evaluation, along with PSNR we report structural similarity wang2004image, mean absolute error (MAE), LPIPS zhang2018unreasonable, and DISTS ding2020image. For clarity, arrows ($\uparrow$ / $\downarrow$) are used in tables and figures to indicate whether higher or lower values correspond to better performance.
TABLE III: Quantitative comparison of different architectures on noisy image reconstruction. SIREN2 (ours) consistently achieves the best or comparable scores across all metrics, showing better noise suppression and structural fidelity.
PSNR($\uparrow$) SSIM($\uparrow$) MAE($\downarrow$) LPIPS($\downarrow$) DISTS($\downarrow$)
2D Riemann (Fig. 11)
Ground truth + Noise 20.48 0.133 0.0734 0.6836 0.3078
SIREN 32.62 0.802 0.0173 0.4139 0.2732
FINER 32.58 0.790 0.0176 0.4397 0.2686
WIRE 28.06 0.449 0.0314 0.6005 0.2864
SIREN2 (Ours) 32.98 0.812 0.0164 0.4096 0.2648
Sparrow
Ground truth + Noise 20.44 0.144 0.0753 0.7077 0.2722
SIREN 29.49 0.905 0.0181 0.4417 0.1816
FINER 31.90 0.880 0.0172 0.4929 0.1423
WIRE 30.23 0.816 0.0205 0.5336 0.1792
SIREN2 (Ours) 32.14 0.883 0.0167 0.4899 0.1414
Nebula
Ground truth + Noise 20.31 0.261 0.0775 0.5554 0.2196
SIREN 25.04 0.664 0.0297 0.5747 0.2419
FINER 26.37 0.697 0.0263 0.5055 0.1503
WIRE 25.02 0.582 0.0350 0.5410 0.1997
SIREN2 (Ours) 26.97 0.695 0.0259 0.4827 0.1120
@@
Table III summarizes the quantitative evaluation of image denoising performance across different models using standard metrics, including PSNR, SSIM, MAE, LPIPS, and DISTS. The results demonstrate that SIREN2 consistently achieves higher fidelity reconstructions with improved perceptual and structural quality compared to the baseline SIREN and other methods. Snapshots of reconstructions presented in Fig. 11 with regard to the 2D-Riemann problem suggest that SIREN and FINER oversmooth fine-scale details, while WIRE preserves global appearance but leaves residual noise. SIREN2 achieves the sharpest and most faithful reconstructions (with the best SSIM and LPIPS), suppressing noise while retaining textures and edges.
@@
VI-D Audio Denoising
TABLE IV: Audio denoising. Best PSNR ($\uparrow$) in dB for different audio clips using FINER++ and SIREN2.
@@ VI-D Audio Denoising
We adopt the same Noise2Self-inspired DIP training procedure from Sec. VI-C for the present audio denoising experiments. Ground-truth audio signal is used only for evaluation. While the general denoising framework is identical, audio signals present distinct challenges. Unlike natural images, which concentrate most energy in low frequencies, audio signals often exhibit a relatively broadband structure (e.g., higher harmonics in music, ambient noise), causing stronger overlap between the underlying signal and Gaussian noise. This overlap makes simple frequency-selective filtering less effective. In an a-priori setting, where no ground-truth information is available, INR-based denoising can provide a useful alternative.
in Sec. VI-A on supervised audio fitting task, $\omega_{0}$ was scaled by a factor of 100 consistently for all network architectures. However, for the present audio-denoising task, a large first-layer periodicity consistently degraded denoising for SIREN, Gauss, WIRE, and FINER, with best denoising $\mathrm{PSNR}<25\,\mathrm{dB}$ for different signals; due to the presence of broadband noise in the noisy signal. We therefore avoid first-layer $\omega_{0}$ scaling for denoising and report results for FINER++ and SIREN2 architectures only which does not use any first layer $\omega_{0}$ scaling. Table IV and Fig. 12 report the accuracy of denoised reconstructions for bach.wav audio clip after 20,000 epochs for FINER++ and SIREN2. The noise scales $s_{0}$ and $s_{1}$ in SIREN2 provide controllable filtering, allowing a principled adjustment of frequency support to match the noise characteristics. For FINER++, varying the bias ranges was evaluated. The results in Table IV show that both networks achieve competitive accuracy in the recovery of the underlying signal.
@@
Figure 12: Audio denoising. Log-spectrogram comparison for audio denoising task performed on bach.wav using FINER++ and SIREN2. SIREN2 achieves higher PSNR and preserves more harmonic detail, as visible in the spectrogram. The rightmost plot shows the PSNR variation of SIREN2 over the $(s_{0},s_{1})$ grid, illustrating that the noise scales $s_{0}$ and $s_{1}$ provide flexibility to control the degree of denoising through modulation of high-frequency suppression.
VI-E Audio Inpainting
We evaluate the inpainting capability of SIREN2 using the same network width, signal length, and noise scales employed in the audio fitting experiments of Sec. VI-A. A fixed percentage of sample values is removed and the model is trained only on the observed subset. Sparsity denotes the retained fraction of samples in the training set. No spatial or structural priors are used; the task reduces to regressing the underlying continuous audio data from incomplete pointwise samples.
Across all sparsity levels, SIREN2 converges to lower reconstruction error for most cases than the baseline SIREN. The gains are consistent for both low and moderate sparsity, indicating that the improved high-frequency receptivity at initialization can benefit recovery of missing regions. The reconstructed samples exhibit sharper transitions indicating influence of initialization scheme on the INR’s ability to approximate the fine-scale details under limited observations.
TABLE V: Audio inpainting. Mean and standard deviation of PSNR (in dB) for different sparsity levels using SIREN and SIREN2.
Audio sample Sparsity SIREN SIREN2
bach.wav 0.05 18.87 $\pm$ 0.28 19.19 $\pm$ 0.03
0.1 22.26 $\pm$ 0.14 22.71 $\pm$ 0.13
0.25 32.79 $\pm$ 0.63 32.97 $\pm$ 0.37
0.5 46.89 $\pm$ 0.33 46.81 $\pm$ 0.44
relay.wav 0.05 16.53 $\pm$ 1.51 18.34 $\pm$ 0.08
0.1 15.92 $\pm$ 0.28 18.10 $\pm$ 0.17
0.25 16.79 $\pm$ 0.07 18.18 $\pm$ 0.18
0.5 17.77 $\pm$ 0.03 19.18 $\pm$ 0.16
tetris.wav 0.05 8.80 $\pm$ 0.13 12.19 $\pm$ 0.02
0.1 9.21 $\pm$ 0.07 11.95 $\pm$ 0.03
0.25 10.13 $\pm$ 0.21 11.87 $\pm$ 0.08
0.5 12.19 $\pm$ 0.27 12.50 $\pm$ 0.08
VII Conclusion
Deep neural networks exhibit a spectral learning bias, where low-frequency components are learned early in training while higher-frequency modes emerge more gradually in the later epochs. However, when the target signal lacks low-frequency content and is dominated by high frequencies, training can suffer from a spectral bottleneck, leading to failure in reconstructing the signal. The present work examined such a phenomenon in the context of sinusoidal representation networks (SIRENs), with a focus on fitting signals dominated by high-frequency content. We analyze the evolution of activation spectra and the neural tangent kernel and characterize the mechanisms underlying the spectral bottleneck. To address this limitation, we propose a target-aware Gaussian weight perturbation scheme, WINNER. The method applies Gaussian noise of a specified scale to the uniformly initialized weights, thereby modifying pre-activation distributions and layerwise spectra. Similar to random Fourier embeddings tancik2020fourier, the present weight perturbation scheme can be used to control the empirical NTK and its eigenbasis, with the added benefit of not introducing additional trainable parameters. The approach achieves state-of-the-art performance on audio fitting tasks and highlights the critical role of initial activation spectra in determining fitting accuracy.
@@ VII Conclusion
Limitations. (a) The proposed initialization scheme relies on prior knowledge or an approximate estimate of the target’s spectral content to determine the perturbation scales $s_{0}$ and $s_{1}$. For problems where the target is not known in advance, such as denoising or solving initial/boundary value PDEs, these hyperparameters can be initialized with estimated values and adaptively updated during early training. A similar limitation also arises when specifying the Fourier scale for other equivalent methods such as random Fourier feature embeddings tancik2020fourier. (b) Similar to the other architectures examined in this study, the training dynamics and final performance of SIREN2 are sensitive to the learning rate and its decay schedule. Without a scheduler, training often results in a strongly oscillatory PSNR evolution resulting in poor convergence. We recommend a decay rate of 1-2% every 20 epochs with an initial learning rate of $10^{-4}$ for both audio and image fitting to achieve stable, non-oscillatory PSNR evolution.
@@ Acknowledgments
The authors HC, DVS, IZ, ZC, and SP gratefully acknowledge the financial assistance provided by Technion - Israel Institute of Technology during the course of present work. HC thanks Sanketh Vedula (Princeton University) for directing him to the work of Sitzmann et al. sitzmann2020implicit; experimenting with it helped the development of noise scheme presented in this work. Generative AI tools were used solely for language editing, limited to grammar and phrasing refinement, and were not used for developing or modifying technical content.
Biographies
Hemanth Chandravamsi received the Ph.D. degree in mechanical engineering from the Technion – Israel Institute of Technology, Haifa, Israel. He is currently a Postdoctoral Researcher with the Faculty of Mechanical Engineering at the Technion. His research interests include scientific and physics aware machine learning, high speed flows, and numerical algorithms.
Dhanush V. Shenoy received the Ph.D. degree in Aerospace Engineering from ISAE-SUPAERO, Toulouse, France, where his doctoral research focused on aeroacoustics and low–Reynolds-number flows. He is currently pursuing post-doctoral research in deep reinforcement learning and its applications to computational fluid dynamics. His research interests include machine learning for flow modeling and control, quantum computing and quantum-inspired algorithms, lattice Boltzmann methods, high-performance computing, and GPU-accelerated numerical solvers.
Itay Zinn received the M.Sc. degree in mechanical engineering from the Ariel University, Israel, and he is currently pursuing his Ph.D. degree in computational fluid dynamics at Technion – Israel Institute of Technology, Israel. His research interests include machine learning, high speed flows, and computational fluid dynamics.
Ziv Chen received the B.Sc. degree from the Technion – Israel Institute of Technology, Haifa, Israel, where he is currently pursuing a direct Ph.D. degree with the Faculty of Electrical and Computer Engineering. He is also affiliated with the Helen Diller Quantum Center, Technion – Israel Institute of Technology, from which he received an excellence scholarship. His research interests include scientific and physics aware machine learning and quantum computing algorithms for physics simulations.
Shimon Pisnoy received the Ph.D. degree in mechanical engineering from the Technion – Israel Institute of Technology, Haifa, Israel, where he is currently a Ph.D. student with the Faculty of Mechanical Engineering. His research interests include machine learning, high speed flows, and computational fluid dynamics.
Steven H. Frankel is a Rosenblatt Chair Professor with the Faculty of Mechanical Engineering, Technion – Israel Institute of Technology, Haifa, Israel. He is also affiliated with the Helen Diller Quantum Center, Technion – Israel Institute of Technology. His research interests include scientific machine learning, computational fluid dynamics, combustion, and quantum computing for physics simulations.